Jessa is 15 years older than her student.In 15 years, twice her age will be thrice the age of her student.Find the present age of her student.
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52798)

.
Let x be the present age of her student.


Then Jessa's present age is (x+15) years.


In 15 years . . . 


    2*((x+15) + 15) = 3*(x+15).


It is your basic equation. Simplify and solve it step by step


    2x + 60 = 3x + 45


    60 - 45 = 3x - 2x  ====>  x = 15.   <<<---=== the student's present age.    ANSWER


CHECK.  Then Jessa's present age is 15 + 15 = 30 years.   ("Jessa is 15 years older than her student.")


        In 15 years: Jessa is 45 years old;  the student is 15 + 15 = 30 years old.


        2*45 = 90 = 3*30.     ! Precisely correct !

Solved.     Congrats  !  !

---------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


Answer by greenestamps(13200)



(1) The difference in their ages will always be 15 years.

(2) In 15 years, twice her age will be thrice the age of her student. In other words, her age will be 3/2 of her student's age.

(3) 3/2 her student's age, with an age difference of 15 years, will be when she is 45 and her student is 30.

(4) Since that will be 15 years from now, she is now 30 and her student is now 15.